28. Compute the specific volume of an air-vapor mixture in cubic meters pre kilogram of dry air when the following conditions prevail : t = 45oC, w = 0.015 kg/kg, and Pt = 100 kPa.


1
Expert's answer
2022-03-14T18:18:05-0400

W = 0.622ps /(pt - ps)

0.015 = 0.622ps /(100 - ps)

ps = 2.355 kPa

pa = pt - ps = 100 - 2.355 = 97.465 kPa

Va = RTa / pa = (0.287)(45+237) / 97.465 = 0.83 m3/kg